OpenStudy (anonymous):

the value of (sin 60)(cos 60) is

OpenStudy (lgbasallote):

are you asked to use trig theorems? or just evaluate it?

OpenStudy (lgbasallote):

coz idk theorems...but i can give you the value

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

\[\sin (60) = \frac{\sqrt 3}{2}\] \[\cos (60) = \frac{1}{2}\] therefore \[(\sin 60)(\cos 60) = \frac{\sqrt 3}{2} \times \frac{1}{2} = \frac{\sqrt 3}{4}\]

OpenStudy (anonymous):

Ty ^^